OpenStudy (anonymous):

I could really use some help, What is p^−m ÷ p^n? p^−m + n p^−m ⋅ n (for sure this isn't it) p^−m − n p^−m ÷ n

OpenStudy (solomonzelman):

The rules of exponents: `Multiplying/Dividing exponents` \(\Large\color{black}{ \displaystyle {\rm \color{red}{a}}^{\rm \color{blue}{b}} \times {\rm \color{red}{a}}^{\rm \color{green}{c}}= {\rm \color{red}{a}}^{{\rm \color{blue}{b}} + {\rm \color{green}{c}}} }\) \(\LARGE\color{white}{ \left| \right| }\) \(\Large\color{black}{ \displaystyle {\rm \color{red}{a}}^{\rm \color{blue}{b}} \div {\rm \color{red}{a}}^{\rm \color{green}{c}}= {\rm \color{red}{a}}^{{\rm \color{blue}{b}} - {\rm \color{green}{c}}} }\) `Negative exponent property.` \(\Large\color{black}{ \displaystyle\left( {\rm \color{red}{D}} \right)^{{\rm \color{green}{-c}}} = \frac{1}{ {\rm \color{red}{D}}^{\color{green}{\rm c}} } }\) \(\Huge\color{white}{ \left|\color{white}{\int^{0}_{3}} \right| }\) those are the basic 3, but if you want all exponential properties, I can suggest you a couple of places: \(\bf\color{green}{{\href{http:///www.mathplanet.com/education/algebra-1/exponents-and-exponential-functions/properties-of-exponents}{mathplanet.com}}}\) or, \(\bf\color{blue}{{\href{http:///hotmath.com/hotmath_help/topics/properties-of-exponents.html}{hotmath.com}}}\) Or if you don't like any of them, then you can google all properties yourself (I apologize, but I'm not planning to make a tutorial in this post right now :P)

OpenStudy (solomonzelman):

(( Only 2 sources, not 3, but ... ))

OpenStudy (solomonzelman):

So you are dividing \(\Large\color{black}{ \displaystyle {\rm \color{red}{p}}^{\rm \color{blue}{-m}} \div {\rm \color{red}{p}}^{\rm \color{green}{n}}= {\rm \color{red}{p}}^{?} }\)

OpenStudy (anonymous):

So if it is mutiplication its going to equal its lesser for example multi=addition and division equals subtraction, So in this case it would be p^−m − n?

OpenStudy (solomonzelman):

Yes, you would get, p^(-m-n)

OpenStudy (anonymous):

Thank you man, I will keep those links saved.

OpenStudy (solomonzelman):

Anytime!